show the equation used for the solution.
In a town election, the winning candidate had 220 more votes than the loser. If 810 votes were cast in all, how many votes did each candidate receive?

Answer by anraymer(35)

T = A + B
T = 810
A = 220 + B


810 = (220 + B) + B
810 = 220 + 2B
590 = 2B
295 votes = B


A = 220 + B
A = 220 + 295
A = 515 votes